# DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition

### v4

@@ 4.1 Framework
Framework Design
We propose the framework to implement two-stage strategy in LLM-based NER. CascadeNER divides NER into two sequential, independently executed, generation-based sub-tasks. In the first sub-task, extraction, the model generate a sentence where all named entities are marked with identifiers and individually re-embeds each entity back into its context, resulting in sentences with identifiers at the number of entities. In the second sub-task, classification, the model receives sentences with identifiers and a list of entity types, and label one entity at a time.
Model Cascading
@@
Pipeline
A simplified pipeline of CascadeNER with an example is shown in Figure 4. Upon receiving the input sentence, CascadeNER first processes the sentence by the extractor to mark all entities with identifiers, and re-embeds each entity back, resulting in sentences with identifiers around the named entities. These sentences are then individually fed into the classifier, which classifies each entity based on the context and the input type list. For multi-granularity data, CascadeNER allows a progressive strategy, significantly improving CasacdeNER’s performance in accurate fine-grained classification. The detailed steps of extraction and classification are discussed in following sections.
@@ 4.2 Extraction
Prompt Design
In the extraction sub-task, we utilize a generation-based extraction method, where special tokens "##" are used to surround any entities identified in the sentence, regardless of the number of entities or their types. For example:
Q: Apple proposes new Macbook
A: ##Apple## proposes new ##Macbook##
This method, compared to conventional sequential extraction, avoids requiring LLMs to perform text alignment, thus reducing task complexity. Comparing similar methods Wang et al. (2023); Hu et al. (2024b), CascadeNER’s query contains only the sentence, without any task descriptions, demonstrations, or category information. The response exclusively uses "##" as the identifiers, and all entities are extracted without specifying categories. CascadeNER achieves low-cost NER by using simple prompts and better generalization by treating all entities uniformly. A detailed comparison with existing methods and further advantages of our method are shown in Appendix F.
Result Fusion
After conducting extensive experiments, we find that the extractor’s precision consistently exceeds recal, regardless of the model or dataset, indicating that while correct entities are effectively identified, there is a tendency for under-detection. To mitigate this issue, we introduce a union strategy in result fusion Ganaie et al. (2022), allowing multiple extraction for one sentence and taking the union of the results to maximize recall. For cases of entity nesting, where different extraction rounds produce overlapping or nested entities, we apply a length-first strategy, retaining the longer entity, as longer entities generally carry more semantic meaning Nguyen and Cao (2010). For example, "Boston University" is semantically more accurate than "Boston" in the context of "She studies in Boston University". The formula of our strategy is shown below:
@@
where $E_{i}$ is the set of extracted entities from the $i$-th extraction, $n$ is the number of extraction rounds, $E_{\text{overlap}}$ is the set of overlapping or nested entities, and $\text{length}(e)$ is the length of entity $e$. The effects of the number of extraction repetitions and other details are provided in Appendix E.1.
@@ 5 Experiment
In this section, we first present the categorization metric changes of DynamicNER before and after dynamic categorization, followed by a comparative analysis of existing methods and CascadeNER’s performance on different versions of DynamicNER. We also conduct experiments of CascadeNER and baselines on existing datasets, along with ablation studies. In evaluations across existing datasets, CascadeNER, with base models fine-tuned using the dynamic version of DynamicNER, demonstrates consistent excellence in all datasets and achieves new SOTA performance in both FewNERD and CrossNER datasets. This confirms that DynamicNER not only provides exceptional effectiveness for evaluating LLM-based NER methods but also offers substantial value in training. Due to space limitations, these experiments are provided in Appendix D and E.
5.1 Categorization Quality Evaluation
To demonstrate that our dynamic categorization improves dataset generalization while maintaining dataset quality, we conduct comparative experiments across three versions of DynamicNER: the Base Version, a version with random parameters for dynamic categorization, and the optimized Dynamic Version. We still employ the 4 metrics for evaluating dataset quality, whose detailed definition are provided in Appendix A. For reliability of the random version’s results, we conduct five independent tests and use the average results. Due to space limitations, we only present results for English and Chinese in Figure 6 here. Other quantitative results are provided in Appendix K.
Language en es fr ru de zh ja kr
# Lists 725 455 501 377 465 786 553 478
5.2 DynamicNER Experiment
Baseline Selection
In our experiments for DynamicNER, we evaluate four NER methods: XLM-RoBERTa Conneau et al. (2020), GPT-NER Wang et al. (2023), PromptNER Ashok and Lipton (2023), and our CascadeNER. XLM-RoBERTa is a famous BERT-based multilingual model widely used as a baseline in multilingual NER research Malmasi et al. (2022); Fetahu et al. (2022), thus being selected as our baseline representing supervised methods. GPT-NER and PromptNER are two major general LLM-based NER methods that achieve performance significantly superior to supervised methods in low-resource scenarios through sophisticated prompt design and powerful GPT models, as discussed in Section 2 and Appendix F.
Model Selection
Given the lack of existing lightweight LLM-based NER methods and controlled variable principles, we evaluate two LLM-based methods and CascadeNER using three LLMs: Qwen2.5-1.5B, Qwen2.5-7B, and GPT-4o. Qwen series is the largest open-source model family. Its lightweight LLMs perform exceptionally across benchmarks and gaining widespread recognition. According to HuggingFace (2024), Qwen2.5-1.5B is the most downloaded open-source model in 2024. Therefore, we select Qwen2.5-1.5B and 7B to represent the current best-performing lightweight LLMs. GPT-4o is the most widely-used current general commercial LLM, and its previous versions are employed in GPT-NER and PromptNER, making it our representative choice for commercial LLMs. In CascadeNER, for statistics convenience, the extractor and the classifier use the same base model.
@@ 5.2 DynamicNER Experiment
Table 3: The results of supervised learning with dynamic version and few-shot learning with dynamic version. G means GPT-NER, P means PromptNER, and C means CascadeNER. The results indicate that, due to its unprecedentedly detailed categorization and multilingual coverage, DynamicNER is a extremely challenging flat NER dataset, placing higher demands on methods’ generalization capability.
Implementation
For the supervised method, XLM-RoBERTa is trained and only evaluated with the base version of DynamicNER. As its fixed classification output layer corresponds to a predefined set of entity types and any modification to the entity type list necessitates full model retraining, it can not be evaluated with the dynamic version. For LLM-based methods, we conduct experiments under three scenarios: supervised learning with base version, supervised learning with dynamic version, and few-shot learning with dynamic version. Training data for GPT-NER and CascadeNER is obtained through format conversion. For PromptNER, as its prompt involves complex designs such as CoT, we utilize LLM-generated prompts by GPT-4o using prompts from its paper as few-shot demonstrations and manually verified the prompts. The repetition count $i$ of CascadeNER for result fusion is set to 3. Potential data contamination are discussed in Appendix H.
@@
Table 4: The results of supervised learning with base version. RoBERT represents XLM-RoBERTa. The supervised method XLM-RoBERTa performs terribly.
Results and Discussion
The results are shown in Table 3 and 4. CascadeNER achieve a significant advantage on DynamicNER, demonstrating its strong generalization and multilingual proficiency. The supervised method XLM-RoBERTa performs terribly, as DynamicNER’s low-resource characteristics make it more suitable for evaluating LLM-based methods. For LLM-based methods, the 3 methods show significant performance variations across different datasets and models. When using GPT-4o and transitioning from supervised to few-shot, PromptNER exhibits notably smaller performance degradation, partially reflecting the generalization advantages of reasoning-focused approaches. However, when migrating to lightweight LLMs, these methods show significantly larger performance drops compared to the other two methods. GPT-NER and CascadeNER demonstrate generally similar performance patterns, but GPT-NER shows more pronounced degradation when migrated to lightweight LLMs, while CascadeNER achieves a greater performance advantage on the dynamic version compared to the base version, validating the effectiveness of the two-stage strategy in both complex classification and reducing computional resource requirement.
@@ 6 Conclusion
This paper introduces DynamicNER, a multi-lingual and multi-granular NER dataset optimized for LLM-based NER, including a human-annotated base version and a dynamic-categorized version. We develop the first dynamic categorization method in NER datasets for DynamicNER, enhancing its generalization while keeping data quality. We also propose CascadeNER, a powerful NER framework which is exceptionally suitable for lightweight LLMs and local deployment, outperforming current LLM-based methods. Moreover, we conduct comprehensive experiments and analyses on DynamicNER and discuss the advantage and future direction of LLM-based NER. More experiments and discussions are provided in Appendix D and F. We hope that DynamicNER and CascadeNER will facilitate future research in LLM-based NER, revitalizing this classical NLP task.
7 Limitation
@@
There are still some challenges for our research. Although CascadeNER is designed to be able to accommodate nested and discontinuous NER, we only conduct evaluation on CascadeNER about flat NER tasks. This limitation arises from the fact that the models in CascadeNER are pre-trained on the dynamic version of DynamicNER, and DynamicNER is a flat NER dataset. Our resources are insufficient to collect enough open-source data for this purpose, which lead to DynamicNER containing only flat NER labels, and thus constraining CascadeNER to flat NER. Furthermore, Due to resource constraints and our failure to find annotators proficient in other languages for manual annotation, DynamicNER currently supports only 8 languages, which somewhat restricts its applicability.
Appendix A Categorization Metric Definition
Cohesion
@@ Appendix A Categorization Metric Definition
Appendix B Categorization Metric Selection
Mixing Types of Different Granularities
@@ Appendix B Categorization Metric Selection
Replace with Synonyms
@@
Remove Irrelevant Types
@@
Merge Types into Miscellaneous
@@
Appendix C Detail of DynamicNER
The specific data volumes for each language are shown in Table 5. It is important to note that for languages except English and Chinese, we partially use manually translated English corpora. This is necessary to balance category distribution, as some languages lack sufficient corpora in specific domains. We also provide conversion scripts that allow DynamicNER to be transformed into train, dev, and test sets with non-overlapping subsets based on coarse categories, making it easier to use for traditional few-shot learning methods.

### v5

@@ 4.1 Framework
Framework Design.
We propose the framework to implement two-stage strategy in LLM-based NER. CascadeNER divides NER into two sequential, independently executed, generation-based sub-tasks. In the first sub-task, extraction, the model generate a sentence where all named entities are marked with identifiers and individually re-embeds each entity back into its context, resulting in sentences with identifiers at the number of entities. In the second sub-task, classification, the model receives sentences with identifiers and a list of entity types, and labels one entity at a time.
Model Cascading.
@@
Pipeline.
A simplified pipeline of CascadeNER is shown in Figure 3. Upon receiving the input sentence, CascadeNER first processes the sentence by the extractor to mark all entities with identifiers, and re-embeds each entity back, resulting in sentences with identifiers around the named entities. These sentences are then individually fed into the classifier, which classifies each entity based on the context and the input type list. For multi-granularity data, CascadeNER allows a progressive strategy, significantly improving CasacdeNER’s performance in accurate fine-grained classification. The detailed steps of extraction and classification are discussed in following sections. Examples about the prompts are provided in Appendix M.
@@ 4.2 Extraction
Prompt Design.
In the extraction sub-task, we utilize a generation-based extraction method, where special tokens "##" are used to surround any entities identified in the sentence, regardless of the number of entities or their types. Figure 5 shows an instance for the prompt and corresponding response.
Figure 5: Example prompts of extraction.
This method, compared to conventional sequential extraction, avoids requiring LLMs to perform text alignment, thus reducing task complexity. Comparing similar methods Wang et al. (2023); Hu et al. (2024b), CascadeNER’s query contains only the sentence, without any task descriptions, demonstrations, or category information. The response exclusively uses "##" as the identifiers, and all entities are extracted without specifying categories. CascadeNER achieves low-cost NER by using simple prompts and better generalization by treating all entities uniformly. A detailed comparison with existing methods and further advantages of our method are shown in Appendix H.
Result Fusion.
After conducting extensive experiments, we find that the extractor’s precision consistently exceeds recall, regardless of the model or dataset, indicating that while correct entities are effectively identified, there is a tendency for under-detection. To mitigate this issue, we introduce a union strategy in result fusion Ganaie et al. (2022), allowing multiple extraction for one sentence and taking the union of the results to maximize recall. For cases of entity nesting, where different extraction rounds produce overlapping or nested entities, we apply a length-first strategy, retaining the longer entity, as longer entities generally carry more semantic meaning Nguyen and Cao (2010). For example, "Boston University" is semantically more accurate than "Boston" in the context of "She studies in Boston University". The formula of our strategy is shown below:
@@
where $E_{i}$ is the set of extracted entities from the $i$-th extraction, $n$ is the number of extraction rounds, $E_{\text{overlap}}$ is the set of overlapping or nested entities, and $\text{length}(e)$ is the length of entity $e$. The effects of the number of extraction repetitions and other details are provided in Appendix F.1.
@@ 5 Experiment
In this section, we first present the categorization metrics of DynamicNER before and after dynamic categorization, followed by a comparative analysis of existing methods and CascadeNER’s performance on different versions of DynamicNER. We also conduct experiments of CascadeNER and baselines on existing datasets in Appendix E, and ablation studies in Appendix F. In evaluations across existing datasets, CascadeNER, with base models fine-tuned using the dynamic version of DynamicNER, demonstrates consistent excellence in all datasets and achieves new SOTA performance in FewNERD and CrossNER. This confirms that DynamicNER not only provides exceptional effectiveness for evaluating LLM-based NER methods but also offers substantial value in training.
@@
Table 2: The results of supervised learning with dynamic version and few-shot learning with dynamic version. G means GPT-NER, P means PromptNER, and C means CascadeNER. The results indicate that, due to its unprecedentedly detailed categorization and multilingual coverage, DynamicNER is an extremely challenging flat NER dataset, placing higher demands on methods’ generalization capability.
5.1 Categorization Quality Evaluation
To demonstrate that our dynamic categorization improves dataset generalization while maintaining dataset quality, we conduct comparative experiments across three versions of DynamicNER: the Base Version, a version with random parameters for dynamic categorization, and the optimized Dynamic Version. We still employ the 4 metrics for evaluating dataset quality, whose detailed definition are provided in Appendix B. For reliability of the random version’s results, we conduct five independent tests and use the average results. Due to space limitations, we only present results for English and Chinese in Figure 6 here. Other quantitative results are provided in Appendix Q.
Language en es fr ru de zh ja kr
# Lists 725 455 501 377 465 786 553 478
5.2 DynamicNER Experiment
Baseline Selection.
In our experiments for DynamicNER, we evaluate four NER methods: XLM-RoBERTa Conneau et al. (2020), GPT-NER Wang et al. (2023), PromptNER Ashok and Lipton (2023), and our CascadeNER. XLM-RoBERTa is a famous BERT-based multilingual model widely used as a baseline in multilingual NER research Malmasi et al. (2022); Fetahu et al. (2022), thus being selected as our baseline representing supervised methods. GPT-NER and PromptNER are two major general LLM-based NER methods that achieve performance significantly superior to supervised methods in low-resource scenarios through sophisticated prompt design and powerful GPT models, as discussed in Section 2 and Appendix H.
Model Selection.
Given the lack of existing lightweight LLM-based NER methods and controlled variable principles, we evaluate two LLM-based methods and CascadeNER using three LLMs: Qwen2.5-1.5B Yang et al. (2024), Qwen2.5-7B, and GPT-4o Hurst et al. (2024). The lightweight LLMs of Qwen series perform exceptionally across benchmarks and gaining widespread recognition. According to HuggingFace (2024), Qwen2.5-1.5B is the most downloaded open-source model in 2024. Therefore, we select Qwen2.5-1.5B and 7B to represent the current best-performing lightweight LLMs. GPT-4o is the most widely-used current general commercial LLM, and its previous versions are employed in GPT-NER and PromptNER, making it our choice. In CascadeNER, the extractor and the classifier use the same base model.
Implementation.
For the supervised method, XLM-RoBERTa is trained and only evaluated with the base version of DynamicNER. As its fixed classification output layer corresponds to a predefined set of entity types and any modification to the entity type list necessitates full model retraining, it cannot be evaluated with the dynamic version. For LLM-based methods, we conduct experiments under three scenarios: supervised learning with base version, supervised learning with dynamic version, and few-shot learning with dynamic version. Training data for GPT-NER and CascadeNER is obtained through format conversion. For PromptNER, as its prompt involves complex designs such as CoT, we utilize LLM-generated prompts by GPT-4o using prompts from its paper as few-shot demonstrations and manually verified the prompts. The repetition count $i$ of CascadeNER for result fusion is set to 3. Potential data contamination are discussed in Appendix N. The finetuning settings and hyperparameters of CascadeNER are provided in Appendix L.
@@ 5.2 DynamicNER Experiment
Table 4: The results of supervised learning with base version. BERT represents XLM-RoBERTa. The supervised method XLM-RoBERTa performs terribly.
Results.
As presented in Table 2 and 4, CascadeNER achieves a significant advantage on DynamicNER, demonstrating its strong generalization. The supervised method XLM-RoBERTa performs terribly, as DynamicNER’s low-resource characteristics make it more suitable for evaluating LLM-based methods. For LLM-based methods, the 3 methods show significant performance variations across different datasets and models. When using GPT-4o and transitioning from supervised to few-shot, PromptNER exhibits notably smaller performance degradation, partially reflecting the generalization advantages of reasoning-focused approaches. However, when migrating to lightweight LLMs, these methods show significantly larger performance drops compared to the other two methods. GPT-NER and CascadeNER demonstrate generally similar performance patterns, but GPT-NER shows more pronounced degradation when migrated to lightweight LLMs, while CascadeNER achieves a greater performance advantage on the dynamic version compared to the base version, validating the effectiveness of the two-stage strategy in the complex classification. CascadeNER also exhibits notable computational efficiency from two key perspectives. First, it is parameter-efficient, leveraging two 7B LLMs to achieve performance superior to baselines that rely on a 100B-level commercial model. Second, its two-stage design enables the use of concise prompts that drastically reduce token consumption compared to competing methods, as discussed in Appendix H, as validated by our cost analysis in Appendix I. This efficiency makes CascadeNER a powerful yet practical solution.
5.3 Performance Across Granularities
To investigate how different methods handle increasing classification complexity, we conduct an analysis across the three granularity levels of DynamicNER. We evaluate all LLM-based methods on the Base version (as the Dynamic version doesn’t have clear granularity levels) for English and Chinese, with results presented in Table 5.
The results lead to two key conclusions. First, as expected, the performance of all models declines as the number of entity types increases from coarse to fine, confirming that fine-grained classification is a significantly more challenging task. More importantly, the analysis reveals a stark difference in how the methods handle this increased complexity. Methods that rely on a single, complex prompt, such as PromptNER, suffer a dramatic performance collapse when scaled to the 155 categories. In contrast, our CascadeNER framework shows remarkable resilience, with a more graceful and controlled degradation in performance. This analysis empirically validates the core motivation for DynamicNER, demonstrating that a challenging, fine-grained benchmark is necessary to expose the limitations of existing methods and accurately evaluate the robustness of purpose-built frameworks.
Model English Chinese
Coarse Medium Fine Coarse Medium Fine
G-1.5B 59.3 55.0 50.2 59.7 54.5 44.1
G-GPT 70.4 69.3 62.4 67.3 65.5 60.4
P-1.5B 58.5 49.2 21.6 57.7 48.8 20.7
P-GPT 75.1 62.5 49.7 73.5 57.1 48.6
C-1.5B 82.7 78.8 67.6 81.2 75.1 64.0
C-GPT 90.6 86.4 77.1 86.5 80.3 74.3
Table 5: F1 scores across granularities on the Base-Supervised setting.
@@ 6 Conclusion
This paper introduces DynamicNER, a multi-lingual and multi-granular NER dataset optimized for LLM-based NER, including a human-annotated base version and a dynamic-categorized version. We develop the first dynamic categorization method in NER datasets for DynamicNER, enhancing its generalization while keeping data quality. We also propose CascadeNER, a powerful NER framework which is exceptionally suitable for lightweight LLMs and local deployment, outperforming current LLM-based methods. Moreover, we conduct comprehensive experiments and analyses on DynamicNER and discuss the advantage and future direction of LLM-based NER. More experiments and discussions are provided in Appendix E and H. We hope that DynamicNER and CascadeNER will facilitate future research in LLM-based NER, revitalizing this classical NLP task.
Limitations
@@
There are still some challenges in our work. Although CascadeNER is designed to be able to accommodate nested and discontinuous NER, we only conduct evaluation on CascadeNER about flat NER tasks. This limitation arises from the fact that the models in CascadeNER are pre-trained on the dynamic version of DynamicNER, and DynamicNER is a flat NER dataset. Our resources are insufficient to collect enough open-source data for this purpose, which lead to DynamicNER containing only flat NER labels, and thus constraining CascadeNER to flat NER. Furthermore, Due to resource constraints and our failure to find annotators proficient in other languages for manual annotation, DynamicNER currently supports only 8 languages, which somewhat restricts its applicability.
Acknowledgements
This work is supported by the National Key R&D Program of China (Grant No. 2024YFC3308304), the "Pioneer" and "Leading Goose" R&D Program of Zhejiang (Grant no. 2025C01128), the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), the State Key Laboratory (SKL) of Biobased Transportation Fuel Technology, and the Zhejiang Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence.
This work is supported in part by the NYUAD Center for Artificial Intelligence and Robotics, funded by Tamkeen under the NYUAD Research Institute Award CG010 and in part by the NYUAD Center for Interdisciplinary Data Science & AI (CIDSAI), funded by Tamkeen under the NYUAD Research Institute Award CG016.
Finally, we are grateful to the three anonymous reviewers for their insightful feedback, which has significantly improved the quality of this manuscript.
Appendix A Dynamic Categorization Strategies
This section elaborates on the four strategies used to transform the manually annotated Base Version of DynamicNER into the final Dynamic Version. This entire process is an algorithm-driven pipeline that systematically modifies the entity type lists and corresponding labels for each data instance to create a more challenging and realistic benchmark. The pipeline consists of four sequential stages. We detail the goal, a concrete example, and the execution process for each strategy below.
Method 1: Mixing Categories of Different Granularities
  • Goal: To test a method’s robustness and generalization capability when faced with entity type lists of varying specificity. This strategy simulates the real-world variability where the required level of detail for an entity can change depending on the context.
  • Example:
    • Sentence: “Ada Lovelace is considered the first computer programmer.”
    • Original State: The entity “Ada Lovelace” is labeled as Person from the coarse-grained list, e.g., [Person, Location, Organization, Product, ...].
    • After Dynamic Change: The system replaces the coarse-grained type with its fine-grained children. The new, shorter type list could be [Artist, Athlete, Scholar, Location, Organization, Product, ...], and the required label for “Ada Lovelace” becomes Scholar.
Method 2: Replacing Types with Synonyms
  • Goal: To evaluate a method’s ability to understand semantic equivalence and not just rely on keyword matching between the entity context and the type list.
  • Example:
    • Sentence: “Bekele won a gold medal in the competition.”
    • Original State: The entity “Bekele” is labeled as Athlete from a type list such as [..., Artist, Author, Athlete, ...].
    • After Dynamic Change: The system replaces the type name Athlete with a valid synonym. The new type list becomes [..., Artist, Author, Sportsperson, ...], and the required label for “Bekele” is now Sportsperson.
Method 3: Removing Irrelevant Categories
  • Goal: To simulate a more realistic, focused NER task where the type list is context-specific, and to test the model’s ability to avoid hallucinating labels that are not in the provided list.
  • Example:
    • Sentence: “In the luxurious suite of The Ritz Carlton Hotel, a traveler is enjoying the beautiful view.”
    • Original State: The entity is “The Ritz Carlton Hotel”, and the type list contains all fine-grained types of Commercial Facility, including many irrelevant ones like Restaurant and Bank.
    • After Dynamic Change: The system prunes unrelated categories. The new type list is [Person, Location, Hotel, Market/Mall, Theater/Cinema, Other Commercial Facility].
Method 4: Merging Types into Miscellaneous/Others
  • Goal: To test the model’s ability to correctly identify and classify infrequent or contextually less important entities into a general catch-all category, a common requirement in practical applications.
  • Example:
    • Sentence: “It was organized by Harry West and constituted a formal electoral pact.”
    • Original State: The entity “Harry West” is labeled as Person from the full type list.
    • After Dynamic Change: The new type list could be [Location, Product, Group, Facility, Art, Miscellaneous], and the required label would become Miscellaneous.
Through this four-stage automated pipeline, DynamicNER generates diverse and challenging evaluation scenarios that go beyond static categorization, enabling a more thorough evaluation of LLMs’ true generalization and contextual understanding capabilities.
Appendix B Categorization Metric Definition
Cohesion.
@@ Appendix B Categorization Metric Definition
Appendix C Categorization Metric Selection
Mixing Types of Different Granularities.
@@ Appendix C Categorization Metric Selection
Replace with Synonyms.
@@
Remove Irrelevant Types.
@@
Merge Types into Miscellaneous.
@@
Appendix D Detail of DynamicNER
The specific data volumes for each language are shown in Table 6. It is important to note that for languages except English and Chinese, we partially use manually translated English corpora. This is necessary to balance category distribution, as some languages lack sufficient corpora in specific domains. For the colloquial portion of the corpora, the Chinese part is sourced from Weibo, while the data for all other languages originated from X platform. We provide conversion scripts that allow DynamicNER to be transformed into train, dev, and test sets with non-overlapping subsets based on coarse categories, making it easier to use for traditional few-shot learning methods.